Use the standard normal probability table to find the value z for each of the following.

(a) P (0≤Z≤z) = 0.41 (b) P(Z≥z) = 0.25

(c) P (Z≤z)= 0.95 (d) P(-z≤Z≤z) = 0.88


1
Expert's answer
2022-05-03T15:05:36-0400

(a) "P (0\u2264Z\u2264z) = 0.41"


"P (Z\u2264z) = 0.41+0.5"

"z=1.340755"

(b) "P (Z\ge z) = 0.25"

"z=0.674490"

(c) "P (Z\u2264z) = 0.95"

"z=1.644854"

(d)"P(-z\u2264Z\u2264z) = 0.88"


"P(Z\ge z) = 0.06"

"z=1.554773"